numerator sqrt5-2x denominator 3 =1

Answer by checkley71(8403)

sqrt(5-2x)/3=1 cross multiply
sqrt(5-2x)=3 square both sides
5-2x=9
-2x=9-5
-2x=4
x=4/-2
x=-2 answer.
proof
sqrt(5-2*-2)/3=1
sqrt(5+4)/3=1
sqrt9/3=1
sqrt9=3
3=3